OpenStudy (anonymous):

Please help ! verify the identity cot(0-pi/2)= -tan 0

OpenStudy (anonymous):

\(cot(0- \frac{\pi}{2}) \) is same as : \(cot(-\frac{\pi}{2})\)

OpenStudy (anonymous):

Now using : \[\cot(-x) = -\cot(x)\] \[\cot(\frac{-\pi}{2}) \implies -\cot(\frac{\pi}{2})\]